OpenStudy (anonymous):

Hi everyone! Can someone help me find a differential operator(annihilator) for x^2cos(4x)? Thanks! :o)

OpenStudy (anonymous):

I haven't run across an example of a x^2 multiplied by a cos4x before. I know separately I would need D^3 for the x^2 and (D^2+16) for the cos4x but how do you do it when they are mashed together like this?

OpenStudy (shadowlegendx):

Pearlz o-o

OpenStudy (shadowlegendx):

Some times I imagine a round white pearly pearl when I read your name...

OpenStudy (perl):

The annhilator for $$ \Large x^{m-1} e^{ax}\cos(bx) ~is~ [(D-a)^2 +b^2]^m $$

OpenStudy (perl):

hi shad

OpenStudy (anonymous):

Really? wow...I ruled that one out because I could find the "alpha" within x^cos4x...where is the alpha?

OpenStudy (anonymous):

zero!

OpenStudy (anonymous):

omg duh! I knew that! :o)

OpenStudy (anonymous):

Quick question Perl...Do you think it's possible to ever come across a function one of the annihilators would work? Like some strange combo?

OpenStudy (anonymous):

no...I was just curious

OpenStudy (anonymous):

is that annihilator you wrote down the same as [D^2-2alphaD + (alpha^2+beta^2)]^n ?

OpenStudy (perl):

yes that is equivalent (note that i used m instead of n )

OpenStudy (anonymous):

okay thanks...I just didn't recognize it...thanks so much! :o)

OpenStudy (perl):

this is a nice list of annhilators

OpenStudy (anonymous):

Hi @Perl, sorry to bother but I have a quick last question... since x^2cos4x has multiplicity of 2 should I choose my annihilator to be: [D^2-2zeroD+(zero^2+beta^2)]^3 Which would simplify to: [D^2+(beta^2)]^3 Is that correct?

OpenStudy (perl):

yes thats correct, the annhilator of the polynomial x^2 will be one degree higher

OpenStudy (anonymous):

so should I expand that out all the way?

OpenStudy (anonymous):

(D^2+16)(D^2+16)(D^2+16)(x^2cos4x)?

OpenStudy (perl):

yes

OpenStudy (anonymous):

wow! lol...okay! :o)

OpenStudy (perl):

but usually we are given a differential equation to solve first

OpenStudy (perl):

P(D)*y = f(x)

OpenStudy (perl):

so you might be given a problem like y '' + 3y ' + y = x^2 cos(x)

OpenStudy (anonymous):

my equation is y"+5y'+6y=x^2cos4x

OpenStudy (anonymous):

yeah I'm good with that procedure thankfully...just having trouble getting this annihilator for some reason...I did fine on all my other work except this one

OpenStudy (anonymous):

look at this>>>

OpenStudy (anonymous):

(D^6+480^4+768^2+4096)(x^2cos4x) doesn't that seem really extreme?

OpenStudy (perl):

$$ \Large { y' '+5y'+6y=x^2\cos(4x) \\ \therefore \\ (D^2 + 5D + 6) y = x^2 cos(4x) \\\\ \text{apply annhilator} \\(D^2 + 16)^3(D^2 + 5D + 6) y =(D^2 + 16)^3 x^2 cos(4x) } $$

OpenStudy (perl):

the right side goes to zero , since it is being annihilated

OpenStudy (anonymous):

I thought you were supposed to expand it out all the way?

OpenStudy (perl):

well we can show that it goes to zero, in steps

OpenStudy (perl):

$$ \Large { (D^2 + 16)(D^2 + 16)(D^2 + 16)(x^2 cos(4x)) \\ = (D^2 + 16)(D^2 + 16)[(D^2\cdot x^2 cos(4x) + 16\cdot x^2 \cos(4x) ] } $$

OpenStudy (perl):

$$ \large { (D^2 + 16)(D^2 + 16)(D^2 + 16)(x^2 cos(4x)) \\ = (D^2 + 16)(D^2 + 16)[(D^2\cdot x^2 cos(4x) + 16\cdot x^2 \cos(4x) ] } $$

OpenStudy (anonymous):

okay...you just jogged my memory...be doing it in steps, we set ourselves up nicely on the left hand side so that we can factor down and find our zeroes

OpenStudy (anonymous):

So when you have to find the "form" of the particular solution, that's where you set all the C1 and C2 etc to A,B,C etc right?

OpenStudy (anonymous):

so yp=A+Bx for example

OpenStudy (perl):

I will show you on my Maple software that it is true that (D^2 + 16)^3 ( x^2 cos(4x)) = 0

OpenStudy (anonymous):

you don't have to...it's okay

OpenStudy (anonymous):

I am working it on my whiteboard right now

OpenStudy (anonymous):

I don't wanna take up all your time

OpenStudy (perl):

$$ \Large{ (D^2 + 16)^3 (x^2 cos(4x)) \\= (D^6+48D^4+768D^2+4096) (x^2 \cos(4x)) } $$

OpenStudy (anonymous):

that's what I have so far...differentiating now

OpenStudy (perl):

its really beautiful on maple: input : f := x^2*cos(4*x); input: diff(f, x$6)+48*diff(f, x$4)+768*diff(f, x$2)+4096*f; output 0

OpenStudy (anonymous):

e-gadz...that looks garbled

OpenStudy (perl):

diff(f, x$6) means the sixth derivative of f , where i defined f above

OpenStudy (anonymous):

Don't get me wrong...I am always interested in seeing new things and discussing theory...I just don't wanna seem like a bother.

OpenStudy (perl):

you really wouldn't do this by hand, it would take forever. annihilator theory ensures you that the right side vanishes (becomes zero)

OpenStudy (anonymous):

OMG! A giant BUG just flew across my ear and I felt it's wings and I screamed and I woke my dog! uhg

OpenStudy (anonymous):

sorry...had to tell someone! :o)

OpenStudy (perl):

:P

OpenStudy (anonymous):

you convinced me...I will stop working this by hand now! :o)

OpenStudy (perl):

i can give you a taste of its complexity

OpenStudy (anonymous):

I just tasted it's complexity...tastes bad!

OpenStudy (anonymous):

bleh :o/

OpenStudy (perl):

$$ \large { D^6 (x^2 cos(4x)) \\ = 7680\cos(4x)-12288x\sin(4x)-4096x^2\cos(4x) \\~\\ D^4 (x^2 cos(4x)) \\= -192\cos(4x)+512x\sin(4x)+256x^2\cos(4x) \\... } $$

OpenStudy (perl):

ok so you want to work on the left side, that part is more interesting

OpenStudy (anonymous):

quick question Perl... So when you have to find the "form" of the particular solution, that's where you set all the C1 and C2 etc to A,B,C etc right? so yp=A+Bx for example

OpenStudy (perl):

correct

OpenStudy (anonymous):

okay thanks!

OpenStudy (perl):

$$ \Large { y' '+5y'+6y=x^2\cos(4x) \\ \therefore \\ (D^2 + 5D + 6) y = x^2 cos(4x) \\\\ \text{apply annhilator} \\(D^2 + 16)^3(D^2 + 5D + 6) y =(D^2 + 16)^3 x^2 cos(4x) \\\therefore \\\\(D^2 + 16)^3(D^2 + 5D + 6) y =0 } $$

OpenStudy (perl):

but its nice you can verify it on a computer algebra system, that the right side 'annihilates'

OpenStudy (perl):

by using the appropriate code or expression

OpenStudy (anonymous):

wow...very neat!

OpenStudy (perl):

this is also a good page on annihilator theory with a good example http://en.wikipedia.org/wiki/Annihilator_method

OpenStudy (anonymous):

omg this left hand side is crazy as well!

OpenStudy (anonymous):

can you confirm the form of my answer?

OpenStudy (anonymous):

I have Yp=Acos4x+Bxcos4x+Cx^2cos4x+Dsin4x+Exsin4x+Fx^2sin4x

OpenStudy (perl):

$$ \Large { y' '+5y'+6y=x^2\cos(4x) \\ \therefore \\ (D^2 + 5D + 6) y = x^2 cos(4x) \\\\ \text{apply annhilator} \\(D^2 + 16)^3(D^2 + 5D + 6) y =(D^2 + 16)^3 x^2 cos(4x) \\\therefore \\\\(D^2 + 16)^3(D^2 + 5D + 6) y =0 \\ roots = 4i, -4i, 4i, -4i, 4i, -4i, -2, -3 \\ y = } $$

OpenStudy (anonymous):

no...they just wanted the form of Yp not Yc also

OpenStudy (perl):

oh ok , i agree then

OpenStudy (anonymous):

whew! that was quite a problem! yikes! Thanks again for all the help! :o) I will try and leave you in peace while I cram for the rest of my exam! Thanks again